# Near-Optimal Decision Trees in a SPLIT Second

### v1

@@ A.9 Appendix Proofs
As the solution to this equation is known to be analytically intractable, we express $u$ in terms of the Lambert $W$ function, which is a well known function that cannot be expressed in terms of elementary functions. Denoted by $W(z)$, this function satisfies the following equation:
@@
A.9.5 Additional Claims, with Proofs
@@
Definition A.4 (Tribes: from (Blanc et al., 2024)).
@@
where $t=\Big{\lfloor}\frac{\ell}{w}\Big{\rfloor}$. (Blanc et al., 2019) prove the following properties of Tribes:
@@
We now use Lemma $7.4$ from (Blanc et al., 2019), which proves the following (again, adjusting for our notation): A random root-to-leaf path of a greedy tree $T$ satisfies the following with probability at least $1-\mathcal{O}(u^{-2})$: If the length of this path is less than $\mathcal{O}(u\log u)$, at any point along that path, all coordinates within the majority block that have not already been queried have correlation at least $\frac{1}{100\sqrt{u}}$. Now, for a greedy tree of depth $d$:
  • We need to set $u\geq\Omega(d\log d)$ so that all root-to-leaf paths have length at most $\mathcal{O}(u\log u)$, so the above lemma applies.
  • Remember that the size of our Tribes block is still fixed as the lookahead depth $d_{l}$, according to Equation 108. From the definition of tribes, all variables in this block will have correlation $\mathcal{O}\Big{(}\frac{\log d_{l}}{d_{l}}\Big{)}$. Because we want the correlations in the majority block to be greater than those in Tribes, we need to set $\frac{1}{100\sqrt{u}}\geq\mathcal{O}\Big{(}\frac{\log d_{l}}{d_{l}}\Big{)}$, implying that $u\leq\mathcal{O}(d_{l}^{2}\log^{2}d_{l})$.
Thus, it follows that $p=\mathcal{O}(u^{-2})$ if the conditions above are satisfied. If we set $d_{l}=\frac{d-1}{2}=\mathcal{O}(d)$, we can say that, for any $\Omega(d\log d)\leq u\leq\mathcal{O}\Big{(}\frac{d^{2}}{\log^{2}}\Big{)}$, a greedy tree of depth $d$ will yield accuracy $\leq\frac{1}{2}+\epsilon$, as it almost never selects any variable from the Tribes block. ∎
@@
  • Sub-block $1$ corresponds to the $d_{l}$ features for which we will compute a parity bit. At a high level, a tree needs to know the parity of the expression in order to ’unlock’ a high accuracy. This also serves to ’trick’ greedy into not choosing these features, because they will have 00 correlation with the label. Let $X_{1}\ldots X_{d_{l}}$ be the features in this sub-block.
@@ A.11 RESPLIT Algorithm
A.12 Indexing Trees in RESPLIT
In this section, we present an algorithm that can quickly index trees output by RESPLIT. Because Algorithm 5 outputs a bespoke data structure involving TreeFARMS objects attached to a set of prefix trees, we needed to devise a method to efficiently query this structure to locate trees at a desired index.
  • For each prefix found by the initial ModifiedTreeFARMS call, we additionally store the number of subtrees that can be formed with that prefix. This is also done for every subsequent node split on beyond the prefix. Algorithm 6 shows how this is done.
  • We also store the cumulative count of the total number of subtrees as we iterate through the list of prefixes. Algorithm 9 called in line $1$ of Algorithm 7 does this.
  • Given a query index, we first perform a binary search over this cumulative count to find the appropriate prefix associated with the queried tree. This is illustrated in lines $2-10$ in Algorithm 7
  • Once the prefix is known, we introduce a variable called tree__\__idx that gives the index of the queried tree relative to that of the prefix, i.e. line $11$ in Algorithm 7. E.g. if we query the $500^{th}$ tree and our prefix has trees indexed $400-600$, the relative index of the query tree within this prefix is $100$.
  • Lastly, using Algorithm 8, we proceed to recursively locate the relevant subtrees beyond the prefix. In particular, at a given node in the prefix, we have access to the number of sub-trees that can be formed with its left and right children. We use this information to create two separate indexes for the left and right child (seen in lines $9-10$)
Algorithm 7 RESPLIT__\__indexing
0:  RESPLIT__\__obj, idx {The RESPLIT object output by Algorithm 5, index of the tree to be found}
1:  $t_{count},p_{counts}$ = RESPLIT__\__Rset__\__Count(RESPLIT__\__obj) {Get ##\## trees in Rashomon set. Call only once.}
2:  $low$, $high$ $=$ $0,len(p_{counts})$
3:  while $low\neq high$ do {Binary Search to find the appropriate prefix tree of the queried tree}
4:     $mid=(low+high)//2$
5:     if $idx<p_{counts}[mid]$ then
6:        $high=mid$
7:     else if $idx>p_{counts}[mid]$ then
8:        $low=mid$
9:     end if
10:  end while
11:  $tree\_idx=idx-p_{counts}[mid]$Subtracting the index by the cumulative tree count of the prefix we’re interested in}
12:  $t_{lookahead}=$RESPLIT__\__obj.prefix__\__list[$mid$]
13:  $tree=$get__\__leaf__\__subtree__\__at__\__idx($t_{lookahead},tree\_idx$) {Use a recursive tree based indexing algorithm to find the appropriate subtrees to fill in for this given prefix tree}
14:  return  $tree$
@@ A.12 Indexing Trees in RESPLIT
Algorithm 9 RESPLIT__\__Rset__\__Count(RESPLIT__\__obj)
0:  RESPLIT__\__obj {The RESPLIT object output by Algorithm 5}
1:  $t_{count}=0$Total ##\## trees}
3:  for $t_{lookahead}\in$ RESPLIT__\__obj do
4:     $p_{count}=1$
5:     for leaf $u\in t_{lookahead}$ do
6:        $tf_{u}=$ TreeFARMS object fitted on subproblem $D(u)$
7:        $s_{count}=len(tf_{u})$Number of subtrees found for subproblem $D(u)$}
8:        $p_{count}=p_{count}\times s_{count}$
9:     end for
10:     $t_{count}=t_{count}+p_{count}$
11:     $p_{counts}.add(t_{count})$
12:  end for
13:  return  $p_{counts},t_{count}$
A.13 Modifications to Existing GOSDT / TreeFARMS Code

### v2

@@ A.8 Appendix Proofs
As the solution to this equation is known to be analytically intractable, we express $u$ in terms of the Lambert $W$ function, which is a well known function that cannot be expressed in terms of elementary functions. Denoted by $W(z)$, the Lambert $W$ function satisfies the following equation:
@@
A.8.5 Additional Claims, with Proofs
@@
Definition A.4 (Tribes: from Blanc et al. 2024).
@@
where $t=\Big{\lfloor}\frac{\ell}{w}\Big{\rfloor}$. Blanc et al. (2019) prove the following properties of Tribes:
@@
We now use Lemma $7.4$ from (Blanc et al., 2019), which proves the following (again, adjusting for our notation): A random root-to-leaf path of a greedy tree $T$ satisfies the following with probability at least $1-\mathcal{O}(u^{-2})$: If the length of this path is less than $\mathcal{O}(\frac{u}{\log u})$, at any point along that path, all coordinates within the majority block that have not already been queried have correlation at least $\frac{1}{100\sqrt{u}}$. Now, for a greedy tree of depth $d$:
  • We need to set $u\geq\Omega(d\log d)$ so that all root-to-leaf paths have length at most $\mathcal{O}\Big{(}\frac{u}{\log u}\Big{)}$, so the above lemma applies.
  • Remember that the size of our Tribes block is still fixed as the lookahead depth $d_{l}$, according to Equation 108. From the definition of tribes, all variables in this block will have correlation $\mathcal{O}\Big{(}\frac{\log d_{l}}{d_{l}}\Big{)}$. Because we want the correlations in the majority block to be greater than those in Tribes, we need to set $\frac{1}{100\sqrt{u}}\geq\Omega\Big{(}\frac{\log d_{l}}{d_{l}}\Big{)}$, implying that $u\leq\mathcal{O}\Big{(}\frac{d_{l}^{2}}{\log^{2}d_{l}}\Big{)}$.
Thus, it follows that $p=\mathcal{O}(u^{-2})$ if the conditions above are satisfied. If we set $d_{l}=\frac{d-1}{2}=\mathcal{O}(d)$, we can say that, for any $\Omega(d\log d)\leq u\leq\mathcal{O}\Big{(}\frac{d^{2}}{\log^{2}d}\Big{)}$, a greedy tree of depth $d$ will yield accuracy $\leq\frac{1}{2}+\epsilon$, as it almost never selects any variable from the Tribes block. ∎
@@
  • Sub-block $1$ corresponds to the $d_{l}$ features for which we will compute a parity bit. At a high level, a tree needs to know the parity of the expression in order to ‘unlock’ a high accuracy. This also serves to ‘trick’ greedy into not choosing these features, because they will have 00 correlation with the label. Let $X_{1}\ldots X_{d_{l}}$ be the features in this sub-block.
@@ A.10 RESPLIT Algorithm
A.11 Indexing Trees in RESPLIT
In this section, we present an algorithm that can quickly index trees output by RESPLIT. This would be especially useful if one wishes to obtain a random sample of trees from the Rashomon set. Because Algorithm 5 outputs a bespoke data structure involving TreeFARMS objects attached to a set of prefix trees, we needed to devise a method to efficiently query this structure to locate trees at a desired index.
  • For each prefix found by the initial ModifiedTreeFARMS call, we additionally store the number of subtrees that can be formed with that prefix. Algorithm 6 shows how this is done.
  • We also store the cumulative count of the total number of trees that can be formed by the prefixes seen so far as we iterate through the list of prefixes. Algorithm 7 called in line $2$ of Algorithm 9 does this.
  • Once the cumulative count is known, we start looping over the entire Rashomon set. For the $i^{th}$ index, we first obtain the corresponding prefix tree and then find the relative index of the $i^{th}$ tree within this prefix tree structure. For example, if we query the $500^{th}$ tree and our prefix contains trees indexed $400-600$ in the Rashomon set, the relative index of the query tree within this prefix is $100$.
  • Using Algorithm 8, we proceed to recursively locate the relevant subtrees beyond the prefix. In particular, at a given node in the prefix, we have access to the number of sub-trees that can be formed with its left and right children. We use this information to create two separate indexes for the left and right child (seen in lines $9-10$)
  • We hash all the indexes for future retrieval (line $16$ in Algorithm 9).
@@ A.11 Indexing Trees in RESPLIT
Algorithm 7 RESPLIT__\__Rset__\__Count(RESPLIT__\__obj)
0:  RESPLIT__\__obj {The RESPLIT object output by Algorithm 5}
1:  $t_{count}=0$Total ##\## trees}
3:  for $t_{lookahead}\in$ RESPLIT__\__obj do
4:     $p_{count}=1$
5:     for leaf $u\in t_{lookahead}$ do
6:        $tf_{u}=$ TreeFARMS object fitted on subproblem $D(u)$
7:        $s_{count}=len(tf_{u})$Number of subtrees found for subproblem $D(u)$}
8:        $p_{count}=p_{count}\times s_{count}$
9:     end for
10:     $t_{count}=t_{count}+p_{count}$
11:     $p_{counts}.add(t_{count})$
12:  end for
13:  return  $p_{counts},t_{count}$
@@
Algorithm 9 RESPLIT__\__indexing
0:  RESPLIT_obj {The RESPLIT object output by Algorithm 5}
[…]
2:  $t_{count},p_{counts}$ = RESPLIT_Rset_Count(RESPLIT_obj) {Total number of trees and prefix-wise cumulative counts}
3:  start = 0
4:  for $i=0$ to $len(p_{counts})-1$ do
5:     if $i>0$ then
6:        start = p_counts[$i-1$]+1 {Start index for prefix $i$}
7:     end if
8:     end = p_counts[$i$] {End index for prefix $i$}
9:     $t_{lookahead}=$ RESPLIT_obj.prefix_list[$i$] {The $i$-th prefix tree}
10:     for local_idx $=0$ to end -- start $-1$ do
11:        global_idx = start + local_idx {Absolute index of tree in Rashomon set}
12:        tree = get_leaf_subtree_at_idx($t_{lookahead},local\_idx$) {Retrieve the corresponding subtree}
13:        hash[global_idx] = tree
14:     end for
15:  end for
16:  return  hash
A.12 Modifications to Existing GOSDT / TreeFARMS Code
@@
A.13 Additional Experimental Results
For thoroughness, we provide several additional experimental results in Figures 20, 21, and 22 and Tables 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, and 17. These experimental results provide other perspectives on the loss/runtime/sparsity tradeoff, with particular emphasis on comparisons with a greedy approach.
Figure 20: Tradeoff between train loss and runtime for all algorithms tested, for different sparsity levels (measured by ##\## of leaves in the tree). Depth Budget $=5$.
Figure 21: Tradeoff between test loss and runtime for all algorithms tested, for different sparsity levels (measured by ##\## of leaves in the tree). Depth Budget $=5$.
Figure 22: Regularized training objective vs. training time (in seconds) for GOSDT vs. our algorithms. The size of the points indicates the number of leaves in the resulting tree. Both SPLIT and LicketySPLIT are much faster for most values of sparsity penalty $\lambda$, with the only potential slowdown being in the sub-second regime due to overhead costs. Depth Budget $=5$.
Table 7: Results for $\lambda=0.001$
Dataset Algorithm Train Objective Runtime (s)
bike GOSDT (SOTA) 0.1322 $\pm$ 0.0015 606.87 $\pm$ 2.12
LicketySPLIT (ours) 0.1328 $\pm$ 0.0010 1.68 $\pm$ 0.05
Greedy 0.2101 $\pm$ 0.0401 0.01 $\pm$ 0.00
adult GOSDT (SOTA) 0.1797 $\pm$ 0.0032 372.62 $\pm$ 2.01
LicketySPLIT (ours) 0.1800 $\pm$ 0.0020 1.46 $\pm$ 0.01
Greedy 0.1950 $\pm$ 0.0350 0.01 $\pm$ 0.00
covertype GOSDT (SOTA) 0.2442 $\pm$ 0.0002 528.67 $\pm$ 1.68
LicketySPLIT (ours) 0.2472 $\pm$ 0.0003 11.86 $\pm$ 0.29
Greedy 0.2681 $\pm$ 0.0020 0.01 $\pm$ 0.00
Table 8: Results for $\lambda=0.006$
Dataset Algorithm Train Objective Runtime (s)
heloc GOSDT 0.3107 $\pm$ 0.0128 266.15 $\pm$ 1.81
SPLIT (ours) 0.3107 $\pm$ 0.0128 5.06 $\pm$ 0.32
LicketySPLIT (ours) 0.3107 $\pm$ 0.0128 0.46 $\pm$ 0.01
compas GOSDT 0.3473 $\pm$ 0.0029 8.99 $\pm$ 0.31
SPLIT (ours) 0.3473 $\pm$ 0.0029 0.64 $\pm$ 0.01
LicketySPLIT (ours) 0.3473 $\pm$ 0.0029 0.29 $\pm$ 0.00
netherlands GOSDT 0.3156 $\pm$ 0.0006 70.49 $\pm$ 0.15
SPLIT (ours) 0.3165 $\pm$ 0.0008 2.29 $\pm$ 0.01
LicketySPLIT (ours) 0.3165 $\pm$ 0.0008 0.61 $\pm$ 0.00
bike GOSDT 0.1736 $\pm$ 0.0007 607.32 $\pm$ 0.89
SPLIT (ours) 0.1737 $\pm$ 0.0072 13.77 $\pm$ 0.18
LicketySPLIT (ours) 0.1753 $\pm$ 0.0158 1.64 $\pm$ 0.01
adult GOSDT 0.1998 $\pm$ 0.0010 264.79 $\pm$ 0.99
SPLIT (ours) 0.1998 $\pm$ 0.0010 4.89 $\pm$ 0.02
LicketySPLIT (ours) 0.1998 $\pm$ 0.0010 1.39 $\pm$ 0.00
covertype GOSDT 0.2652 $\pm$ 0.0001 32.27 $\pm$ 0.27
SPLIT (ours) 0.2652 $\pm$ 0.0001 6.21 $\pm$ 0.02
LicketySPLIT (ours) 0.2652 $\pm$ 0.0001 11.24 $\pm$ 0.01
Table 9: Results for $\lambda=0.011$
Dataset Algorithm Train Objective Runtime (s)
heloc GOSDT 0.3214 $\pm$ 0.0017 231.41 $\pm$ 5.55
SPLIT (ours) 0.3214 $\pm$ 0.0017 4.10 $\pm$ 0.03
LicketySPLIT (ours) 0.3214 $\pm$ 0.0017 0.44 $\pm$ 0.00
compas GOSDT 0.3621 $\pm$ 0.0066 7.17 $\pm$ 0.53
SPLIT (ours) 0.3621 $\pm$ 0.0066 0.32 $\pm$ 0.08
LicketySPLIT (ours) 0.3621 $\pm$ 0.0066 0.26 $\pm$ 0.01
netherlands GOSDT 0.3406 $\pm$ 0.0006 56.54 $\pm$ 0.11
SPLIT (ours) 0.3515 $\pm$ 0.0105 1.74 $\pm$ 0.01
LicketySPLIT (ours) 0.3515 $\pm$ 0.0105 0.60 $\pm$ 0.00
bike GOSDT 0.1961 $\pm$ 0.0006 610.20 $\pm$ 0.55
SPLIT (ours) 0.1961 $\pm$ 0.0006 12.91 $\pm$ 0.05
LicketySPLIT (ours) 0.1961 $\pm$ 0.0006 1.60 $\pm$ 0.01
adult GOSDT 0.2148 $\pm$ 0.0010 211.86 $\pm$ 0.68
SPLIT (ours) 0.2148 $\pm$ 0.0010 2.90 $\pm$ 0.07
LicketySPLIT (ours) 0.2148 $\pm$ 0.0010 1.39 $\pm$ 0.00
covertype GOSDT 0.2752 $\pm$ 0.0001 2.62 $\pm$ 0.23
SPLIT (ours) 0.2752 $\pm$ 0.0001 6.07 $\pm$ 0.08
LicketySPLIT (ours) 0.2752 $\pm$ 0.0001 10.70 $\pm$ 0.00
Dataset Binarization Time (s) Algorithm Runtimes (s) Test Loss
compas [2.69, 2.91] LicketySPLIT (ours) [2.85, 2.97] [0.306, 0.328]
SPLIT (ours) [2.84, 3.01] [0.314, 0.335]
CART [0.00, 0.00] [0.322, 0.354]
bank [0.37, 0.41] LicketySPLIT (ours) [0.47, 0.52] [0.103, 0.119]
SPLIT (ours) [0.57, 0.67] [0.101, 0.116]
CART [0.00, 0.00] [0.106, 0.123]
bike [2.20, 2.32] LicketySPLIT (ours) [2.67, 2.81] [0.133, 0.147]
SPLIT (ours) [3.80, 4.71] [0.139, 0.152]
CART [0.01, 0.01] [0.207, 0.215]
adult [2.26, 2.74] LicketySPLIT (ours) [2.94, 3.19] [0.155, 0.159]
SPLIT (ours) [3.66, 4.26] [0.155, 0.159]
CART [0.02, 0.02] [0.183, 0.191]
hypothyroid [0.99, 1.32] LicketySPLIT (ours) [1.13, 1.30] [0.004, 0.005]
SPLIT (ours) [1.18, 1.36] [0.004, 0.005]
CART [0.00, 0.00] [0.009, 0.014]
covertype [19.49, 20.00] LicketySPLIT (ours) [25.50, 25.75] [0.242, 0.244]
SPLIT (ours) [25.62, 25.91] [0.242, 0.244]
CART [0.68, 0.69] [0.266, 0.269]
netherlands [1.92, 2.04] LicketySPLIT (ours) [2.31, 2.39] [0.285, 0.294]
SPLIT (ours) [2.74, 3.06] [0.284, 0.294]
CART [0.00, 0.00] [0.314, 0.338]
heloc [0.89, 1.01] LicketySPLIT (ours) [1.20, 1.27] [0.284, 0.289]
SPLIT (ours) [2.08, 2.53] [0.284, 0.289]
CART [0.01, 0.01] [0.293, 0.299]
spambase [0.70, 0.73] LicketySPLIT (ours) [0.88, 0.90] [0.094, 0.114]
SPLIT (ours) [1.16, 1.19] [0.097, 0.111]
CART [0.01, 0.01] [0.164, 0.207]
Table 10: Results (##\## leaves between 3–6). The $95\%$ confidence interval is shown. Binarization is only applicable to LicketySPLIT/SPLIT. The runtimes for SPLIT / LicketySPLIT include binarization time.
Dataset Binarization Time (s) Algorithm Runtimes (s) Test Loss
compas [2.69, 2.91] LicketySPLIT (ours) [2.86, 2.98] [0.303, 0.321]
SPLIT (ours) [2.91, 3.06] [0.302, 0.320]
CART [0.00, 0.00] [0.325, 0.338]
bank [0.37, 0.41] LicketySPLIT (ours) [0.48, 0.53] [0.103, 0.118]
SPLIT (ours) [0.59, 0.68] [0.101, 0.117]
CART [0.00, 0.00] [0.106, 0.123]
bike [2.20, 2.32] LicketySPLIT (ours) [2.70, 2.84] [0.123, 0.125]
SPLIT (ours) [4.06, 4.78] [0.127, 0.134]
CART [0.01, 0.01] [0.166, 0.238]
adult [2.26, 2.74] LicketySPLIT (ours) [2.97, 3.22] [0.149, 0.154]
SPLIT (ours) [3.95, 4.54] [0.149, 0.155]
CART [0.03, 0.04] [0.165, 0.180]
hypothyroid [0.99, 1.32] LicketySPLIT (ours) [1.13, 1.30] [0.003, 0.005]
SPLIT (ours) [1.20, 1.38] [0.003, 0.005]
CART [0.00, 0.00] [0.002, 0.004]
covertype [19.49, 20.00] LicketySPLIT (ours) [25.50, 25.74] [0.240, 0.243]
SPLIT (ours) [26.57, 26.92] [0.239, 0.242]
CART [0.99, 1.00] [0.254, 0.256]
netherlands [1.92, 2.04] LicketySPLIT (ours) [2.31, 2.41] [0.282, 0.293]
SPLIT (ours) [2.79, 3.12] [0.282, 0.293]
CART [0.00, 0.00] [0.297, 0.314]
heloc [0.89, 1.01] LicketySPLIT (ours) [1.21, 1.27] [0.284, 0.293]
SPLIT (ours) [2.18, 2.42] [0.282, 0.293]
CART [0.00, 0.00] [0.291, 0.327]
spambase [0.70, 0.73] LicketySPLIT (ours) [0.89, 0.91] [0.085, 0.096]
SPLIT (ours) [1.36, 1.52] [0.085, 0.098]
CART [0.02, 0.02] [0.114, 0.141]
Table 11: Results (##\## leaves between 7–10). The $95\%$ confidence interval is shown. Binarization is only applicable to LicketySPLIT/SPLIT. The runtimes for SPLIT / LicketySPLIT include binarization time.
Dataset Binarization Time (s) Algorithm Runtimes (s) Test Loss
compas [2.69, 2.91] SPLIT (ours) [2.92, 3.08] [0.302, 0.316]
CART [0.00, 0.00] [0.318, 0.333]
bank [0.37, 0.41] LicketySPLIT (ours) [0.49, 0.53] [0.099, 0.116]
SPLIT (ours) [0.59, 0.69] [0.100, 0.118]
CART [0.00, 0.00] [0.104, 0.119]
bike [2.20, 2.32] LicketySPLIT (ours) [2.70, 2.83] [0.114, 0.123]
SPLIT (ours) [4.40, 5.14] [0.121, 0.129]
CART [0.02, 0.02] [0.130, 0.139]
adult [2.26, 2.74] LicketySPLIT (ours) [2.97, 3.22] [0.148, 0.155]
SPLIT (ours) [4.09, 4.79] [0.148, 0.154]
CART [0.04, 0.04] [0.154, 0.161]
netherlands [1.92, 2.04] LicketySPLIT (ours) [2.32, 2.42] [0.283, 0.291]
SPLIT (ours) [2.89, 3.22] [0.282, 0.291]
CART [0.00, 0.00] [0.293, 0.309]
heloc [0.89, 1.01] LicketySPLIT (ours) [1.23, 1.29] [0.281, 0.292]
SPLIT (ours) [2.41, 2.73] [0.286, 0.297]
CART [0.02, 0.02] [0.298, 0.306]
spambase [0.70, 0.73] LicketySPLIT (ours) [0.89, 0.92] [0.086, 0.094]
SPLIT (ours) [1.50, 1.65] [0.081, 0.093]
CART [0.02, 0.02] [0.114, 0.136]
Table 12: Results (##\## leaves between 11–14). The $95\%$ confidence interval is shown. Binarization is only applicable to LicketySPLIT/SPLIT. The runtimes for SPLIT / LicketySPLIT include binarization time.
Dataset Binarization Time (s) Algorithm # Leaves (95% CI) Runtimes (s) Test Loss
compas [2.69, 2.91] LicketySPLIT (ours) [14.2, 17.2] [2.99, 3.23] [0.305, 0.325]
SPLIT (ours) [13.6, 16.0] [2.76, 3.32] [0.304, 0.314]
CART [28.6, 30.8] [0.00, 0.00] [0.307, 0.324]
bank [0.37, 0.41] LicketySPLIT (ours) [20.2, 23.8] [0.52, 0.57] [0.102, 0.116]
SPLIT (ours) [15.2, 16.0] [0.63, 0.75] [0.102, 0.118]
CART [16.6, 17.8] [0.01, 0.01] [0.098, 0.112]
bike [2.20, 2.32] LicketySPLIT (ours) [17.8, 19.6] [2.72, 2.94] [0.114, 0.122]
CART [27.6, 28.2] [0.02, 0.03] [0.122, 0.130]
adult [2.26, 2.74] LicketySPLIT (ours) [19.8, 21.4] [2.90, 3.36] [0.146, 0.151]
SPLIT (ours) [15.2, 16.0] [4.29, 5.45] [0.148, 0.153]
CART [22.6, 23.2] [0.02, 0.03] [0.152, 0.157]
netherlands [1.92, 2.04] LicketySPLIT (ours) [15.4, 18.4] [2.33, 2.72] [0.282, 0.291]
SPLIT (ours) [13.4, 15.2] [2.97, 3.45] [0.284, 0.291]
CART [18.6, 20.2] [0.01, 0.01] [0.293, 0.306]
heloc [0.89, 1.01] LicketySPLIT (ours) [18.0, 23.0] [1.25, 1.44] [0.285, 0.295]
SPLIT (ours) [14.4, 16.4] [2.31, 2.63] [0.286, 0.301]
CART [21.4, 23.2] [0.02, 0.02] [0.290, 0.299]
spambase [0.70, 0.73] LicketySPLIT (ours) [24.4, 25.6] [0.91, 0.96] [0.081, 0.088]
SPLIT (ours) [14.0, 15.8] [1.46, 1.57] [0.081, 0.093]
CART [20.0, 23.2] [0.02, 0.02] [0.082, 0.103]
Table 13: Comparing CART and SPLIT/LicketySPLIT for non sparse trees. The $95\%$ confidence interval is shown. Binarization is only applicable to LicketySPLIT/SPLIT. The runtimes for SPLIT / LicketySPLIT include binarization time. We report the best tree with between $15-30$ leaves found during hyperparameter search.
Dataset Leaves Runtimes Losses
bank [20.20, 23.80] [0.52, 0.56] [0.102, 0.116]
bike [17.80, 19.60] [2.80, 2.92] [0.114, 0.122]
adult [19.80, 21.40] [3.18, 3.40] [0.146, 0.151]
netherlands [15.40, 18.40] [2.34, 2.43] [0.282, 0.292]
heloc [18.00, 23.00] [1.24, 1.30] [0.286, 0.295]
spambase [24.40, 25.60] [0.88, 0.91] [0.081, 0.088]
Table 14: SPLIT/LicketySPLIT for non-sparse trees. For this variant, we set $\lambda=1e-5$ and ran our algorithms over $5$ trials. We show the $95\%$ confidence interval. This shows that our algorithms are capable of producing non-sparse trees. We may not prefer to do this in practice if there are interpretability constraints or if we can get a well performing model with much fewer than $20$ leaves.
Dataset Algorithm Runtimes (s) Test Loss
compas CART (with binary features) [0.00, 0.00] [0.325, 0.352]
CART (with cont features) [0.00, 0.00] [0.322, 0.354]
bank CART (with binary features) [0.00, 0.00] [0.106, 0.122]
CART (with cont features) [0.00, 0.00] [0.106, 0.123]
bike CART (with binary features) [0.02, 0.03] [0.208, 0.215]
CART (with cont features) [0.01, 0.01] [0.208, 0.215]
adult CART (with binary features) [0.01, 0.01] [0.168, 0.197]
CART (with cont features) [0.02, 0.02] [0.183, 0.191]
hypothyroid CART (with binary features) [0.00, 0.00] [0.009, 0.014]
CART (with cont features) [0.00, 0.00] [0.009, 0.014]
covertype CART (with binary features) [0.06, 0.07] [0.253, 0.256]
CART (with cont features) [0.68, 0.69] [0.266, 0.269]
netherlands CART (with binary features) [0.01, 0.01] [0.332, 0.342]
CART (with cont features) [0.00, 0.00] [0.314, 0.338]
heloc CART (with binary features) [0.01, 0.01] [0.293, 0.299]
CART (with cont features) [0.01, 0.01] [0.293, 0.299]
spambase CART (with binary features) [0.00, 0.00] [0.156, 0.208]
CART (with cont features) [0.01, 0.01] [0.164, 0.207]
Table 15: Comparison between CART (with binary features) and CART (with cont features) for trees with 3–6 leaves. The $95\%$ confidence interval is shown.
Dataset Algorithm Runtimes (s) Test Loss
compas CART (with binary features) [0.00, 0.00] [0.3197, 0.3388]
CART (with cont features) [0.00, 0.00] [0.3253, 0.3385]
bank CART (with binary features) [0.00, 0.00] [0.1109, 0.1193]
CART (with cont features) [0.00, 0.00] [0.1061, 0.1227]
bike CART (with binary features) [0.01, 0.01] [0.1718, 0.1911]
CART (with cont features) [0.00, 0.01] [0.1661, 0.2380]
adult CART (with binary features) [0.02, 0.02] [0.1647, 0.1803]
CART (with cont features) [0.03, 0.04] [0.1647, 0.1803]
hypothyroid CART (with binary features) [0.00, 0.00] [0.0030, 0.0053]
CART (with cont features) [0.00, 0.00] [0.0015, 0.0042]
covertype CART (with binary features) [0.07, 0.07] [0.2501, 0.2557]
CART (with cont features) [0.99, 1.00] [0.2540, 0.2560]
netherlands CART (with binary features) [0.01, 0.01] [0.3323, 0.3590]
CART (with cont features) [0.00, 0.00] [0.2966, 0.3137]
heloc CART (with binary features) [0.00, 0.00] [0.2895, 0.3015]
CART (with cont features) [0.01, 0.01] [0.2926, 0.2990]
spambase CART (with binary features) [0.00, 0.00] [0.1101, 0.1407]
CART (with cont features) [0.02, 0.02] [0.1142, 0.1409]
Table 16: Comparison between CART (with binary features) and CART (with cont features) for trees with 7–10 leaves. The $95\%$ confidence interval is shown.
Dataset Algorithm Runtimes (s) Test Loss
compas CART (with binary features) [0.00, 0.00] [0.3179, 0.3334]
CART (with cont features) [0.00, 0.00] [0.3176, 0.3334]
bank CART (with binary features) [0.00, 0.00] [0.1087, 0.1176]
CART (with cont features) [0.00, 0.00] [0.1045, 0.1193]
bike CART (with binary features) [0.04, 0.05] [0.1288, 0.1354]
CART (with cont features) [0.02, 0.02] [0.1296, 0.1387]
adult CART (with binary features) [0.02, 0.03] [0.1543, 0.1615]
CART (with cont features) [0.04, 0.04] [0.1542, 0.1615]
hypothyroid CART (with binary features) [0.00, 0.00] [0.0045, 0.0083]
CART (with cont features) [0.00, 0.00] [0.0023, 0.0045]
covertype CART (with binary features) [0.07, 0.07] [0.2485, 0.2537]
CART (with cont features) [1.22, 1.23] [0.2529, 0.2554]
netherlands CART (with binary features) [0.01, 0.01] [0.3021, 0.3115]
CART (with cont features) [0.00, 0.00] [0.2934, 0.3090]
heloc CART (with binary features) [0.01, 0.01] [0.2961, 0.3047]
CART (with cont features) [0.02, 0.02] [0.2983, 0.3057]
spambase CART (with binary features) [0.01, 0.01] [0.1064, 0.1416]
CART (with cont features) [0.02, 0.02] [0.1144, 0.1359]
Table 17: Comparison between CART (with binary features) and CART (with cont features) for trees with 11–14 leaves. The $95\%$ confidence interval is shown.
A.14 Predictive Multiplicity of our Rashomon Set
We illustrate another metric showing the approximation ability of RESPLIT. For each example in the training set, we computed the variance in predictions across models in the Rashomon set. The distribution of this variance over training examples is shown as a box plot for each dataset. Figure 23 and Table 18 shows that there is minimal empirical difference in the predictive multiplicity of original vs RESPLIT Rashomon sets.
Figure 23: Illustration of the predictive multiplicity of the original and RESPLIT Rashomon Sets. $\lambda=0.02$, $\epsilon=0.01$, Depth Budget $=5$, Lookahead depth $=3$.
Dataset Mean Variance (Original Rashomon Set) $\pm$ Std Dev Mean Variance (RESPLIT) $\pm$ Std Dev
bike $0.2464\pm 0.0023$ $0.2458\pm 0.0024$
netherlands $0.2017\pm 0.0190$ $0.2187\pm 0.0122$
hiv $0.2485\pm 0.0018$ $0.2478\pm 0.0027$
compas $0.2389\pm 0.0138$ $0.2407\pm 0.0149$
heloc $0.2419\pm 0.0081$ $0.2419\pm 0.0081$
spambase $0.2359\pm 0.0055$ $0.2284\pm 0.0037$
Table 18: Illustration of the predictive multiplicity of the original and RESPLIT Rashomon Sets. $\lambda=0.02$, $\epsilon=0.01$, Depth Budget $=5$, Lookahead depth $=3$. This presents results in Figure $1$ in tabular form.